# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ 3.2 Generalized communication sliding
$\langle v,\widehat{u}^{T}-u\rangle+\phi(\widehat{u}^{T})-\phi(u)$ $\leq\frac{2\eta}{T(T+3)}D_{w}(u,x^{init})+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}({u},x_{i})$
$\quad-\frac{(T+1)(T+2)}{T(T+3)}\eta D_{w}(u,u^{T})-\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(\widehat{u}^{T},x_{i})+\frac{4M^{2}}{\eta(T+3)}.$
$\langle v,\widehat{u}^{T}-u\rangle+\phi(\widehat{u}^{T})-\phi(u)$ $\leq\sum_{i\in\mathcal{I}}\eta_{i}D_{w}({u},x_{i})-\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(\widehat{u}^{T},x_{i})$
$\quad-(\frac{\mu}{C}+\eta)D_{w}(u,u^{T})+\frac{2M^{2}/\eta}{T(T+1)}\sum_{t=1}^{T}\frac{\lambda_{t}}{\beta_{t}},$
3.3 Convergence of multi-timescale gradient sliding
The proof of convergence of Algorithm 1 follows a similar type of argument as the proof of convergence of PDHG and DCS: the primal updates (19) and dual updates control the following two terms ((C), (C)):
+$\qquad+\left\{\sum_{i_{s}=0}^{N_{s}-1}\langle-K_{s}\widetilde{X}_{s}^{i_{s}},y_{s}^{i_{s}}-y_{s}\rangle+r_{s}(R^{*}_{s}(y_{s}^{i_{s}})-R^{*}_{s}(y_{s}))\right\}.$
The above sum (approximately) matches the gap $\sum_{k=0}^{N}Q(\widehat{Z}^{k},Z)$ up to an additive term ((C))
$\sum_{s=1}^{S}\sum_{k=0}^{N}\langle\widehat{X}^{k}-\widetilde{X}_{s}^{\lfloor k/r_{s}\rfloor},K^{*}_{s,v}(y_{s}-\overline{y}_{s}^{k})\rangle=\sum_{i_{s}=0}^{N_{s}-1}\langle\sum_{i=0}^{r_{s}-1}\widehat{x}_{v}^{r_{s}i_{s}+i}-\widetilde{x}_{s,v}^{i_{s}},K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle.$
Notice that as compared to the proof of PDHG and DCS, due to the different timescales for the duals, we bound the above terms at dual time scales: instead of controlling $\langle\widehat{x}_{v}^{k}-\widetilde{x}_{s,v}^{\lfloor k/r_{s}\rfloor},K^{*}_{s,v}(y_{s}-y_{s}^{\lfloor k/r_{s}\rfloor})\rangle$ for each $k$, we control the cumulative term (sum from $k=r_{s}i_{s}$ to $k=r_{s}(i_{1}+1)-1$). With our choice of the $\widetilde{X}_{s}^{i_{s}}$ and the mixture terms used in primal proximal updates, the result follows. We defer the proof to Appendix C.
Lemma 3.1.
Under Assumption 2.4 and the first part of Assumption 2.3, with the following choice of parameters:
  • $\alpha_{s,i_{s}}=\alpha=1$, $\theta_{k}=1$, $T_{k}=T\geq 1$;
  • $\eta_{k,s}=\eta\rho_{s}$ where $\rho_{s}\geq 0$ and $\sum_{s=1}^{S}\rho_{s}=1$;
  • $\tau_{s}=\frac{2\widetilde{\kappa}_{s}^{2}}{\rho_{s}\eta}$ where $\widetilde{\kappa}_{s}:=\sup_{\|y_{s}\|\leq 1}\|K_{s}^{*}y_{s}\|_{*}$;
  • $\lambda_{t}=t+1$ and $\beta_{t}=t/2$ for the CS procedure for all iteration $k$.
Then for any $Z\in\overline{\mathcal{X}}\times\mathbb{R}^{n}$,
$\quad(N+1)\cdot Q(Z^{N};Z)$
$\leq\eta\left\{\frac{3}{2}(\sum_{s=1}^{S}r_{s}\rho_{s})D_{w^{X}}(X,X^{init})-D_{w^{X}}(X,X^{N})\right\}$
$\quad+\frac{1}{\eta}\left\{\sum_{s=1}^{S}\frac{\widetilde{\kappa}_{s}^{2}r_{s}}{\rho_{s}}\left\{3D_{w^{y}_{s}}(y_{s},y_{s}^{init})-D_{w^{y}_{s}}(y_{s},y_{s}^{N_{s}-1})\right\}+\frac{4mM^{2}(N+1)}{T+3}\right\}$
Thus, with proper choices of $\eta,\rho_{s},T$, we get the following bound.
Theorem 3.1.
$D_{w^{X}}(\widehat{X},X^{init})\leq D^{X}<\infty,\quad\sup_{y_{s}\in\mathsf{dom}(R^{*}_{s})}D_{w^{y}_{s}}(y_{s},y_{s}^{init})\leq D_{s}^{y}<\infty.$
Under the conditions in Lemma 3.1, taking $\eta=(\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}})\sqrt{\frac{8}{3D^{X}}}$, $\rho_{s}=\frac{\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}}}{\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}}}$, and $T\geq\lfloor\frac{4mM^{2}(N+1)}{\overline{r}(\sum_{s=1}^{S}\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}})^{2}}\rfloor$ where $\overline{r}:=\sum_{s=1}^{S}r_{s}\rho_{s}$, we have
$\sup_{Y^{\prime}\in\mathbb{R}^{n}}Q(Z^{N};\widehat{X},Y^{\prime})\leq\frac{2\sqrt{6}\overline{r}\cdot(\sum_{s=1}^{S}\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}})\cdot\sqrt{D^{X}}}{N+1}.$ (22)
Proof of Theorem 3.1.
From Lemma 3.1, we first notice that with $\rho_{s}=\frac{\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}}}{\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}}}$
Thus, we have
$T+3\geq\frac{4mM^{2}(N+1)}{\overline{r}(\sum_{s=1}^{S}\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}})^{2}}\implies\frac{4mM^{2}(N+1)}{T+3}\leq\sum_{s=1}^{S}\frac{\widetilde{\kappa}_{s}^{2}r_{s}D_{s}^{y}}{\rho_{s}}$
Thus, with the additional assumptions, we get
$\sup_{Y^{\prime}\in\mathbb{R}^{n}}Q(Z^{N};\widehat{X},Y^{\prime})$ $\leq(N+1)^{-1}\left\{\frac{3\eta D^{X}}{2}(\sum_{s=1}^{S}r_{s}\rho_{s})+\frac{4}{\eta}(\sum_{s=1}^{S}\frac{\widetilde{\kappa}_{s}^{2}r_{s}D_{s}^{y}}{\rho_{s}})\right\}$
We would like to point out that Algorithm 1 and the above guarantees Lemma 3.1 and Theorem 3.1 (as well as Lemma 4.1 and Theorem 4.1 below) hold for any saddle point problem of the structure ($\mathcal{P}_{s}$), where the duals are block-separable. In particular, the convergence holds without the assumptions specific to distributed optimization problems, such as the assumption that $\ker(K)=\mathsf{Span}(\mathbf{1})$ (Assumption 2.2), or $R_{s}$ has to be nonnegative and $R_{s}^{-1}(\mathbf{0})=\{\mathbf{0}\}$ (second part of Assumption 2.3). In addition, Lemma 3.1 (and Lemma 4.1 below) can be used to show the convergence of $\sup_{Z\in\overline{\mathcal{Z}}}Q(Z^{N};Z)$, with $D^{X}\geq\sup_{X\in\overline{\mathcal{X}}}D_{w^{X}}(X,X^{init})$. In Theorem 3.1 (and Theorem 4.1), we give the “weaker” convergence for a fixed $\widehat{X}$. This is due to our Lemmas 2.2, 2.3, and Corollary 2.1, which only require an upper bound on $\sup_{Y^{\prime}\in\mathbb{R}^{n}}Q(Z;X^{*},Y^{\prime})$. In addition, as will be seen in Section 4.1, the dependence on $D_{w^{X}}(X^{*},X^{init})$ rather than $\sup_{X\in\overline{\mathcal{X}}}D_{w^{X}}(X,X^{init})$ is crucial to obtain complexities which depend on function similarities for AMT-GS.
In addition, we point out that according to Theorme 3.1, $T=O(\frac{mM^{2}N}{\overline{r}(\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}})^{2}})$, and so the total subgradient needed to find an $\epsilon$ suboptimal solution is
$NT=O(\frac{mM^{2}N^{2}}{\overline{r}(\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}})^{2}})=O(\frac{\overline{r}mM^{2}D^{X}}{\epsilon^{2}}).$
3.4 Communication round complexities for ($\mathcal{P}$)
With additional assumptions specific to distributed optimization, and with proper choices of $R_{s}$’s, the duality gap for ($\mathcal{P}_{s}$) can be related to the suboptimality in terms of objective values $F$ and/or violation of the consensus constraint for the original problem ($\mathcal{P}$). Next, we establish such connection.
We point out that with $w_{s}^{y}=\frac{1}{2}\|y_{s}\|_{2}^{2}$ and $y_{s}^{0}=\mathbf{0}$, we can take $\sqrt{2D_{s}^{y}}$ as $\frac{\xi+a_{s}}{\sigma_{\min}^{+}(K_{s})}$ for ${R}_{s}^{ccv}$ in (5), as $\frac{(1+\xi)a_{s}}{\sigma_{\min}^{+}(K_{s})}$ for ${R}_{s}^{prj}$ in (6), as $\frac{\xi+\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$ for $\widehat{R}_{s}^{ccv}$ in (12), and as $\frac{(1+\xi)\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$ for $\widehat{R}_{s}^{prj}$ in (13).
Corollary 3.2.
@@ 3.4 Communication round complexities for ($\mathcal{P}$)
Assume that all norms are the $l_{2}$ norm, and take $y_{s}^{0}=\mathbf{0}$, $w_{s}^{y}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$ and $w^{x}(x)=\frac{1}{2}\|x\|^{2}$. Assume that Assumptions 2.1, 2.2 and 2.3 and the conditions of Theorem 3.1 hold, and $\{f_{v}\}_{v\in V}$ is $\widehat{a}_{1}$-similar. Take $\rho_{s}=\frac{\|K_{s}\|}{\sum_{s^{\prime}=1}^{S}\|K_{s^{\prime}}\|}$, $\overline{r}=\sum_{s=1}^{S}r_{s}\rho_{s}$, and assume $D_{w^{X}}(X^{*},X^{init})\leq D^{X}$ , then for
$N\geq\frac{2\sqrt{3}\overline{r}A\sqrt{D^{X}}}{\epsilon},$
@@
    $\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=\widehat{R}_{s}^{ccv}$ as defined in (12) and $A=\frac{(\sum_{s=1}^{S}\|K_{s}\|)}{\sigma_{\min}^{+}(K)}\cdot(\xi+\widehat{a}_{1})$;
@@
    $\Pi(\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=\widehat{R}_{s}^{prj}$ as defined in (13) and $A=\frac{(1+\xi)(\sum_{s=1}^{S}\|K_{s}\|)\cdot\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$.
Corollary 3.3.
Assume that all norms are the $l_{2}$ norm, and take $y_{s}^{0}=\mathbf{0}$, $w_{s}^{y}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$ and $w^{x}(x)=\frac{1}{2}\|x\|^{2}$. Assume that $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$ for all $s,s^{\prime}\in[S]$, that Assumptions 2.1, 2.2 and 2.3 and the conditions of Theorem 3.1 hold, and that $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar. Take $\overline{r}=\sum_{s=1}^{S}r_{s}\rho_{s}$, and assume $D_{w^{X}}(X^{*},X^{init})\leq D^{X}$ , then for
$N\geq\frac{2\sqrt{3}\overline{r}A\sqrt{D^{X}}}{\epsilon},$
  1. 1.
    $\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=R_{s}^{ccv}$ as defined in (5), $\rho_{s}=(\frac{\xi+a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{\xi+a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})})$, and $A=\sum_{s=1}^{S}(\xi+a_{s})\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})}$;
  2. 2.
    $\Pi(\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=R_{s}^{prj}$ satisfies (6), $\rho_{s}=(\frac{a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})})$, and $A=(1+\xi)(\sum_{s=1}^{S}a_{s}\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})})$.
@@
3.5 Discussions
Below, we look at the communication rounds ($N$) and the subgradient oracle complexities ($T$ and $NT$) of Algorithm 1 in order to find an $\epsilon$ suboptimal solution. In Section 3.5.1, we focus on the saddle point formulation ($\mathcal{P}_{s}$), assuming that the regularization $R$ is given. In Section 3.5.2, we look at the original distributed optimization problem ($\mathcal{P}$), and choose $R$ based on the discussion in Section 2.3.
3.5.1 Saddle point problem ($\mathcal{P}_{s}$)
Theorem 3.1 implies that to find an $\epsilon$-suboptimal solution, one can take
$N=O(\frac{\overline{r}\cdot(\sum_{s=1}^{S}\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}})\cdot\sqrt{D^{X}}}{\epsilon}),\quad T=O(\frac{M^{2}m\sqrt{D^{X}}}{\epsilon\sum_{s=1}^{S}\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}}}),$
where $\overline{r}:=\sum_{s=1}^{S}r_{s}\rho_{s}=\frac{\sum_{s=1}^{S}r_{s}\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}}}{\sum_{s=1}^{S}\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}}}$. In particular, the total rounds of communication is $O(\frac{\overline{r}}{\epsilon})$. For the subgradient, notice that $NT=O(\frac{\overline{r}M^{2}mD^{X}}{\epsilon^{2}})$, which agrees with [lan_communication-efficient_2020] for when $r_{s}=1$ for all $s$.
@@ 3.5.1 Saddle point problem ($\mathcal{P}_{s}$)
Then, with the above $N$, the dual variable $y_{s}$ is updated $O(\frac{\overline{r}}{r_{s}\epsilon})$ times, which is different for duals with different $r_{s}$. Thus, suppose one is allowed to choose the update frequencies $\{r_{s}\}_{s\in[S]}$, to minimize the total cost to find an $\epsilon$ suboptimal solution, the following should be (approximately) minimized
[…]
$\sum_{s\in\mathcal{S}}c_{s}\leq(\sum_{l=0}^{|\mathcal{S}|-1}\xi^{-l})\cdot\max_{s\in\mathcal{S}}c_{s}\leq\frac{\xi}{\xi-1}\cdot\max_{s\in\mathcal{S}}c_{s}\leq\frac{\xi}{\xi-1}\cdot\sum_{s\in\mathcal{S}}c_{s},$
and so the additive cost is a constant approximation to the maximum.
3.5.2 Distributed optimization problem ($\mathcal{P}$)
For the distributed optimization problem ($\mathcal{P}$), the more natural measure of performance is the objective value suboptimality and the consensus constraint violation, as defined in (2). As indicated by Corollaries 3.2 and 3.3, with good choices of the regularization $R$, $\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k}$ or its projection using $\Pi$ satisfies (2). In the rest of Section 3.5.2, we assume that all norms are $l_{2}$ norms, and take $w^{x}(x)=\frac{1}{2}\|x\|^{2}$ and $w^{y}_{s}(x)=\frac{1}{2}\|y_{s}\|^{2}$.
Bounds using the Lipschitz constants. Consider the case where $\|f^{\prime}_{v}\|\leq M_{f}$, and to guarantee that $\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k}$ is an $(\epsilon,\epsilon/\xi)$-solution, in Corollary 3.2 we can take $\xi=\widehat{a}_{1}=2\sqrt{m}M_{f}$ which gives the following $N_{1}$, and in Corollary 3.3, we can take $\xi=a_{s}=2\sqrt{m}M_{f}$ for all $s\in[S]$, which gives the following $N_{2}$:
Both $N_{1}$ and $N_{2}$ depend linearly in $\overline{r}$. However, when $\{K_{s}\}_{s\in[S]}$ are orthogonal, as discussed in Section 2.3, $\sigma_{\min}^{+}(K)\leq\sigma_{\min}^{+}(K_{s})$ for all $s$, and so in terms of the rounds of communication $N$, it appears that orthogonality allows a more refined (i.e. $s$-dependent) control over the decomposition of the function variation and thus the dual domain size, thereby achieving better convergence. In addition, similar to the argument in Section 3.5.1, when the cost of updating $y_{s}$ is $c_{s}$ and total cost is additive, one should choose $r_{s}\propto\sqrt{c_{s}/\|K_{s}\|}$ when Corollary 3.2 holds, and $r_{s}\propto\sqrt{c_{s}/(\|K_{s}\|/\sigma_{\min}^{+}(K_{s}))}$ if Corollary 3.3 holds. Similar results hold for $\Pi(\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k})$ to be an $(\epsilon,0)$-solution.
Bounds using the function similarity. In reality, sometimes the functions $\{f_{v}\}_{v\in V}$ exhibit similarity. For instance, in the extreme case $f_{v}=f_{v^{\prime}}$ for all $v,v^{\prime}\in V$, and thus communication is not needed at all! In that case, the bound on $\sup_{\widetilde{X}\in\overline{\mathcal{X}},~{}K\widetilde{X}=\mathbf{0}}\|(I-\Pi)\nabla F(\widetilde{X})\|$ (and other terms using $\Pi_{s}$) using the Lipschitz constant $M_{f}$ is too loose: in fact, one can choose $\widehat{a}_{1}=a_{s}=\epsilon_{0}$ for all $s$ for arbitrarily small $\epsilon_{0}>0$, then when $R_{s}$ are set according to (13) or (6) with constant $\xi$, one only needs $N=O(\frac{\epsilon_{0}}{\epsilon}\sum_{s=1}^{S}r_{s})$, which can be arbitrarily small.
More generally, choosing $\xi=1$ and setting $R_{s}$ according to (13) or (6), we obtain the following bound on the rounds of communication following Corollary 3.2 ($N_{3}$) and 3.3 ($N_{4}$)
@@ 3.5.2 Distributed optimization problem ($\mathcal{P}$)
In fact, when $S=1$, such dependency is optimal. Indeed, [ArjevaniShamir2015Communication] designs a pair of “chain like” functions $\{F_{1},F_{2}\}$, such that for any $\gamma\geq 0$, $\{\gamma F_{1},\gamma F_{2}\}$ is $\sqrt{1.5}\gamma$-similar. In addition, when $m/2$ agents are given $\gamma F_{1}$ and the rest are give $\gamma F_{2}$, finding an $\epsilon$ suboptimal $x$ (in terms of the objective value) in the $l_{2}$ unit ball requires $\Omega(\frac{\gamma}{\epsilon/m})$ rounds of communication (see Theorem 2 and the discussions after it in [ArjevaniShamir2015Communication]). For our algorithm, with $\widehat{a}_{1}=O(\sqrt{m}\gamma)$, $D^{x}=1/2$, and $K=I-\Pi$ (and so $\|K\|=\sigma_{\min}^{+}(K)$), we have $N_{3}=O(\frac{r_{1}\gamma}{\epsilon/m})$. Thus, $y_{1}$ is updated only $N_{3}/r_{1}=O(\frac{\gamma}{\epsilon/m})$ times, which is also the number of actual communication rounds needed. This achieves the theoretical lower bound, and so is optimal.
However, to achieve the above function-variation-dependent bounds, the parameters $\widehat{a}_{1}$ and $\{a_{s}\}_{s\in[S]}$ need to be set correctly. It is an interesting open question how one can achieve such dependence without additional prior knowledge (such as function similarity) about $\{f_{v}\}_{v\in V}$.
@@
Thus, from the cost-minimization perspective in Section 3.5.1, denoting the cost of updating $y_{s}$ as $c_{s}$, one should choose $r_{s}\propto\sqrt{\frac{c_{s}/|\mathsf{Des}(s)|}{\sqrt{\sup_{x\in\mathcal{X}}\mathsf{Var}_{i\sim\mu_{s}}(\overline{f^{\prime}}_{i}(x))}}}$. This corroborates the intuition that if along some $K_{s}$ the function does not vary by too much ($\mathsf{Var}_{i\sim\mu_{s}}(\overline{f^{\prime}}_{i})$ is small), then $\mathsf{Agent}(y_{s})$ does not need to update $y_{s}$ very frequently (can use larger $r_{s}$).
4 Accelerated convergence under strong convexity
The convergence result in Section 3.3 can be applied to objectives $\{f_{v}\}_{v\in V}$ where $\mu=0$. In case strong convexity holds, i.e. $\mu>0$, Algorithm 1 can achieve the accelerated convergence rate $O(1/N^{2})$. We defer the proof of Lemma 4.1 to Appendix D.

### v2

@@ 4.5 Convergence for general convex objectives
Assume that all norms are the $l_{2}$ norm, and take $y_{s}^{0}=\mathbf{0}$, $w_{y_{s}}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$, $w_{x}(x)=\frac{1}{2}\|x\|^{2}$, and $w_{X}(X)=\sum_{v\in V}w_{x}(x_{v})$. Assume that conditions of Corollary 3.3 hold, $D(X^{*},X^{init})\leq D^{X}$, and the following holds for $A$ specified below
$N\geq\frac{2\sqrt{3}\overline{r}A\sqrt{D^{X}}}{\epsilon}.$
@@
    If $\{f_{v}\}_{v\in V}$ is $\widehat{a}_{1}$-similar, take $\rho_{s}=\frac{\|K_{s}\|}{\sum_{s^{\prime}=1}^{S}\|K_{s^{\prime}}\|}$
    1. (a)
      $\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=\widehat{R}_{s}^{ccv}$ as defined in (27) (then $\sqrt{2D_{s}^{y}}=\frac{\xi+\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$) and $A=\frac{(\sum_{s=1}^{S}\|K_{s}\|)}{\sigma_{\min}^{+}(K)}\cdot(\xi+\widehat{a}_{1})$;
    2. (b)
      $\Pi(\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=\widehat{R}_{s}^{prj}$ as defined in (28) (then $\sqrt{2D_{s}^{y}}=\frac{(1+\xi)\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$) and $A=\frac{(1+\xi)(\sum_{s=1}^{S}\|K_{s}\|)\cdot\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$.
@@
    If $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar,
    1. (a)
      $\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=R_{s}^{ccv}$ as defined in (20) (then $\sqrt{2D_{s}^{y}}=\frac{\xi+a_{s}}{\sigma_{\min}^{+}(K_{s})}$), $\rho_{s}=(\frac{\xi+a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{\xi+a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})})$, and $A=\sum_{s=1}^{S}(\xi+a_{s})\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})}$;
    2. (b)
      $\Pi(\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=R_{s}^{prj}$ satisfies (21) (then $\sqrt{2D_{s}^{y}}=\frac{(1+\xi)a_{s}}{\sigma_{\min}^{+}(K_{s})}$), $\rho_{s}=(\frac{a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})})$, and $A=(1+\xi)(\sum_{s=1}^{S}a_{s}\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})})$.
@@
Bounds using the Lipschitz constants. Consider the case where $\|f^{\prime}_{v}\|\leq M_{f}$, and to guarantee that $\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k}$ is an $(\epsilon,\epsilon/\xi)$-solution, in Corollary 4.2, with $\{f_{v}\}_{v\in V}$ $\widehat{a}_{1}$-similar, we can take $\xi=\widehat{a}_{1}=2\sqrt{m}M_{f}$ which gives the following $N_{1}$, and with $\{f_{v}\}_{v\in V}$ $\{(a_{s},K_{s})\}_{s\in[S]}$-similar, we can take $\xi=a_{s}=2\sqrt{m}M_{f}$ for all $s\in[S]$, which gives the following $N_{2}$:
Both $N_{1}$ and $N_{2}$ depend linearly in $\overline{r}$. However, when $\{K_{s}\}_{s\in[S]}$ are orthogonal, as discussed in Section 4.3, $\sigma_{\min}^{+}(K)\leq\sigma_{\min}^{+}(K_{s})$ for all $s$, and so in terms of the rounds of communication $N$, it appears that orthogonality allows a more refined (i.e. $s$-dependent) control over the decomposition of the function variation and thus the dual domain size, thereby achieving better convergence. In addition, similar to the argument in Section 3.2, when the cost of updating $y_{s}$ is $c_{s}$ and total cost is additive, one should choose $r_{s}\propto\sqrt{c_{s}/\|K_{s}\|}$ under $\widehat{a}_{1}$-similarity, and $r_{s}\propto\sqrt{c_{s}/(\|K_{s}\|/\sigma_{\min}^{+}(K_{s}))}$ under $\{(a_{s},K_{s})\}_{s\in[S]}$-similarity. Similar results hold for $\Pi(\frac{1}{N+1}\sum_{k=0}^{N}\widehat{X}^{k})$ to be an $(\epsilon,0)$-solution.
@@
More generally, choosing $\xi=1$ and setting $R_{s}$ according to (28) or (21), we obtain the following bound on the rounds of communication following under $\widehat{a}_{1}$-similarity ($N_{3}$) and $\{(a_{s},K_{s})\}_{s\in[S]}$-similarity ($N_{4}$):
@@
In fact, when $S=1$, such dependency is optimal. Indeed, [4] designs a pair of “chain like” functions $\{F_{1},F_{2}\}$, such that for any $\gamma\geq 0$, $\{\gamma F_{1},\gamma F_{2}\}$ is $\sqrt{1.5}\gamma$-similar. In addition, when $m/2$ agents are given $\gamma F_{1}$ and the rest are give $\gamma F_{2}$, finding an $\epsilon$ suboptimal $x$ (in terms of the objective value) in the $l_{2}$ unit ball requires $\Omega(\frac{\gamma}{\epsilon/m})$ rounds of communication (see Theorem 2 and the discussions after it in [4]). For our algorithm, with $\widehat{a}_{1}=O(\sqrt{m}\gamma)$, $D^{x}=1/2$, and $K=I-\Pi$ (and so $\|K\|=\sigma_{\min}^{+}(K)$), we have $N_{3}=O(\frac{r_{1}\gamma}{\epsilon/m})$. Thus, $y_{1}$ is updated only $N_{3}/r_{1}=O(\frac{\gamma}{\epsilon/m})$ times, which is also the number of actual communication rounds needed. This achieves the theoretical lower bound, and so is optimal.
@@
Thus, from the cost-minimization perspective in the discussion in Section 3.2, denoting the cost of updating $y_{s}$ as $c_{s}$, one should choose $r_{s}\propto\sqrt{\frac{c_{s}/|\mathsf{Des}(s)|}{\sqrt{\sup_{x\in\mathcal{X}}\mathsf{Var}_{i\sim\mu_{s}}(\overline{f^{\prime}}_{i}(x))}}}$. This corroborates the intuition that if along some $K_{s}$ the function does not vary by too much ($\mathsf{Var}_{i\sim\mu_{s}}(\overline{f^{\prime}}_{i})$ is small), then $\mathsf{Agent}(y_{s})$ does not need to update $y_{s}$ very frequently (can use larger $r_{s}$).
4.6 Convergence for strongly convex objectives
Good initialization for ($\mathcal{P}_{d}^{lift}$). In Corollary 3.4, assuming that $\mathcal{X}$ is compact, then one can always use $D^{X}\geq\sup_{X\in{\mathcal{X}}}D_{w^{X}}(X,X^{init})$, suggesting that $X^{init}$ should be chosen as the “center” of $\mathcal{X}$}, and $D^{X}$ measures the (squared) radius of $\mathcal{X}$}. The resulting $N$, then, depends on $D^{X}$. However, such dependence on the size of $\mathcal{X}$} could be suboptimal, especially when local objectives are similar. Indeed, in the extreme case where all local functions are the same, then primal agents can optimize their local objectives without communication at all.
To take advantage of potential similarities in the local functions, we propose initializing the primal variables at (approximate) local optimal solutions, which has the following guarantee on $D(X^{*},X^{init})$.